OpenStudy (anonymous):

Find the sum of the series. ∑(n = 0, ∞) 2/9^n

OpenStudy (anonymous):

\[\sum_{n=0}^\infty \left(\frac{2}{9}\right)^n~~\text{or}~~\sum_{n=0}^\infty\frac{2}{9^n}~~?\]

OpenStudy (anonymous):

the second!

OpenStudy (anonymous):

Recall the formula: \[\sum_{n=0}^\infty ar^n=\frac{a}{1-r}\] You can write your series to match this formula to easily evaluate the sum: \[\sum_{n=0}^\infty 2\left(\frac{1}{9}\right)^n=\frac{2}{1-\frac{1}{9}}=\frac{9}{4}\]

OpenStudy (anonymous):

This is very helpful, thank you!!

OpenStudy (anonymous):

yw